OpenStudy (anonymous):

Simplify (3√ 7) + (5√ 7)

OpenStudy (zepp):

Treat \(\sqrt{7}\) as x And factor: \((3+5)\sqrt{7}\)

OpenStudy (anonymous):

5/4?

OpenStudy (zepp):

It's like if you had \(3x+5x=(3+5)x=8x\)

OpenStudy (anonymous):

8/7

OpenStudy (zepp):

\((8)\sqrt{7}\)

OpenStudy (anonymous):

so 8/7 is right

OpenStudy (zepp):

No

OpenStudy (zepp):

\(8\sqrt{7}\), not \(\large\frac{8}{7}\)

OpenStudy (anonymous):

o 8 squareroot 7

OpenStudy (anonymous):

so i had it righ t.. moving on to next question